OpenStudy (anonymous):

find the equation of a curve that passes thru (1,3)and has slope of y/x^2 at any point of x/y f(x) = sqrt (x) + 3; g(x) = (x+3)/2

OpenStudy (dumbcow):

what are f and g for ?

OpenStudy (dumbcow):

anyway looks like you set slope equal to y/x^2 \[\frac{dy}{dx} = \frac{y}{x^{2}}\] separate variables and integrate both sides \[\int\limits_{ }^{} \frac{dy}{y} = \int\limits_{}^{} \frac{dx}{x^{2}}\] \[\ln y = -\frac{1}{x} + C\] take e raised to both sides to cancel log \[y = k e^{-1/x}\] where k is constant e^C plug in point (1,3) to find k --> k = 3e \[\large y = 3e^{1-\frac{1}{x}}\]